# Iterative Compositional Data Generation for Robot Control

### v2

@@ Appendix B Additional Experimental Details
This section provides details of the experimental setting used to obtain all results in Section 4 in the main paper.
@@ B.1.1 Diffusion Model Training
We train diffusion models on transition tuples $(s,a,r,s^{\prime},d)$ with total dimension $164$: a 77-dimensional state, an 8-dimensional action, a scalar reward, a 77-dimensional next state, and a binary terminal indicator. All continuous dimensions are standardized (zero mean, unit variance) using statistics computed from the pooled dataset of all training tasks. The terminal indicator remains unnormalized and is discretized with a threshold of $0.5$.
All model variants use the same Elucidated Diffusion schedule; differences arise only from the denoiser architecture and a small number of optimization hyperparameters.
@@ B.3 Task List
This yields $4\times 4\times 4=64$ tasks. We generate a random split using seed 0, producing 32 training and 32 test tasks with no overlap. We use the first 14 training tasks for all diffusion and multitask policy experiments, and evaluate on all 32 held-out test tasks. Table 6 lists the tasks, and Figures 10 and 11 visualize the split.

### v3

@@ Appendix A Regime to Study Compositionality in CompoSuite
Figure 10: Performance of the monolithic SynthER-based diffusion architecture and the semantic compositional diffusion transformer over iterations of the self-improvement procedure in the $56/256$-task training regime. (Left) Zero-shot best success rate achieved so far. RL agents trained on synthetic data from the semantic compositional diffusion architecture consistently achieve higher success rates than agents trained on data from the monolithic SynthER-style diffusion architecture. (Middle) Number of tasks solved at least once across iterations. The semantic compositional diffusion transformer enables policies to solve a substantially larger fraction of tasks at least once compared to the monolithic baseline. (Right) Zero-shot best success rate achieved so far, separated by initial task difficulty. Tasks are partitioned into those that exhibit non-zero success at iteration 0 and those that were entirely unsolved. The semantic compositional architecture particularly improves performance on tasks for which some initial success is already obtained. Shaded regions indicate standard error over 33 diffusion seeds.
To further validate that this regime exhibits behavior consistent with our IIWA-only setting, we repeat the iterative data generation procedure in the $56/256$-task regime and directly compare the monolithic SynthER-style diffusion model with our semantic compositional diffusion transformer architecture. Figure 10 reports the evolution of zero-shot success rates and solved-task coverage over iterations of Algorithm 1.
We observe similar trends to those in Figure 6: both architectures benefit from iterative self-improvement, while policies trained on data generated by the semantic compositional model consistently achieve higher zero-shot success rates and solve a larger fraction of tasks at least once across iterations. Notably, the performance gap between the monolithic and semantic compositional model is slightly reduced in the $56/256$ setting. This is expected, as increasing the number of compositional axes while preserving a comparable training-task ratio exposes the monolithic model to a broader set of task combinations, partially alleviating the severity of the generalization challenge. Importantly, the compositional model continues to exhibit a clear advantage, indicating that the benefits of exploiting task structure persist beyond the IIWA-only setting.
Finally, separating tasks by initial difficulty reveals a similar asymmetry to that observed in the main experiments. While the semantic compositional model improves in both regimes, it exhibits substantially larger gains on tasks that achieve non-zero success at iteration 0, with more gradual improvements on initially unsolved tasks. This behavior is consistent with the interpretation that the model refines factor-specific encoder–decoder representations using self-generated data. Overall, these results support the conclusion that the $14/64$ IIWA-only setting constitutes a particularly challenging low-data regime, and that the qualitative self-improvement dynamics of the semantic compositional diffusion model are stable as the number of compositional task axes increases.
@@ Appendix B Additional Experimental Details
This section provides details of the experimental setting used to obtain all results in the paper.
@@ B.1.1 Diffusion Model Training
We train diffusion models on transition tuples $(s,a,r,s^{\prime},d)$ with total dimension $164$: a 77-dimensional state, an 8-dimensional action, a scalar reward, a 77-dimensional next state, and a binary terminal indicator. All continuous dimensions are standardized (zero mean, unit variance) using statistics computed from the pooled dataset of all training tasks. The terminal indicator remains unnormalized and is discretized with a threshold of $0.5$. All model variants use the same Elucidated Diffusion schedule; differences arise only from the denoiser architecture and a small number of optimization hyperparameters.
@@ B.3 Task List
This yields $4\times 4\times 4=64$ tasks. We generate a random split using seed 0, producing 32 training and 32 test tasks with no overlap. We use the first 14 training tasks for all diffusion and multitask policy experiments, and evaluate on all 32 held-out test tasks. Table 6 lists the tasks, and Figures 11 and 12 visualize the split.